OpenStudy (anonymous):

A new drug has been found to be effective in treating 75% of the people afflicted by a certain disease. If the drug is administered to 400 people who have this disease, what are the mean and the standard deviation of the number of people for whom the drug can be expected to be effective? (Round your standard deviation to two decimal places.)

OpenStudy (anonymous):

mean ??? people standard deviation ???? people

OpenStudy (btaylor):

I don't think you have enough information.

OpenStudy (kropot72):

This can be solved by using the normal approximation to the binomial distribution. Let n be the number of people who have the disease and have the drug administered. Let p be the probability that the drug is effective. Then the mean number of people for whom the drug is effective is given by: \[np=400\times0.75=you\ can\ calculate\] The standard deviation is given by: \[\sqrt{np(1-p)}=\sqrt{400\times0.75\times0.25}=you\ can\ calculate\]